OpenStudy (anonymous):

What is the slope of the line 2x – 5y = -10?

OpenStudy (anonymous):

is it 5/2?

OpenStudy (blurbendy):

-5y = -2x - 10 y = 2/5x + 2 slope = 2/5

OpenStudy (anonymous):

thankyou

OpenStudy (blurbendy):

yw

OpenStudy (anonymous):

can you help me with another @blurbendy